What's idempotent law?. Please explain. Urgent .exams tomorrow!

Grade:11

3 Answers

Ajay
7 years ago

The defintion is easily availible on google,  may be you should try that more frequently.

Idempotent Laws:

For any finite set A;

(i) A U A = A

(ii) A ∩ A = A

Umakant biswal
7 years ago
@ riya
Idempotent Law
. In Boolean algebra, Idempotent Law states that combining a quantity with itself either by logical addition or logical multiplication will result in a logical sum or product that is the equivalent of the quantity. 
example – A + A = A; A x A = A
hope it clears your doubt 
all the best 

 
jyoti bhatia
7 years ago
In Boolean algebra, Idempotent Law states that combining a quantity with itself either by logical addition or logical multiplication will result in a logical sum or product that is the equivalent of the quantity (A+A = A; A x A = A).
Idempotent Law is a law which states that an element x of an algebraic system satisfies x^2 = x.
 In a Boolean algebra, both of the dyadic operations are idempotent.